1.00 mL of a 3.50 × 10–4 M solution of oleic acid is diluted with 9.00 mL of petroleum ether, forming solution A. 2.00 mL of solution A is diluted with 8.00 mL of petroleum ether, forming solution B. How many grams of oleic acid are 5.00 mL of solution B? (molar mass for oleic acid = 282 g/mol)




1
Expert's answer
2021-04-15T06:00:59-0400

Volume"=10ml"

Solution A"=" "3.5\times 10^{-4}\over 0.001" "=3.5\times 10^{-7}M"

Solution B

Volume"=10ml"

"n=3.5\times 10^{-7}\times 0.2=7.0\times 10^{-8}"

Moles of Oleic acid in "5ml" solution.

"B=7.0\times 10^{-8}\times 0.5=3.5\times 10^{-8}"

Mass of oleic acid

"3.5\times 10^{-8}\times 282=9.8\times 10^{-6}g"